# Explainable Procedural Mistake Detection

### v1

@@ Acknowledgments
This work was supported in part by the DARPA PTG program HR00112220003. We would like to thank Megan Su, Ruixuan Deng, Fengyuan Hu, Andy Chung, as well as the entire MSRP team for their helpful discussions, feedback, and insights.
@@ Appendix A Ego4D-PMD Data Curation Details
To alleviate these challenges and focus our inquiry, we follow Du et al. (2023) in recasting Ego4D, a procedural video dataset with a breadth of narrated everyday actions, into a mistake detection format. The collection of the Ego4D Grauman et al. (2022) for Procedural Mistake Detection (Ego4D-PMD) data consisted of several preprocessing steps, outlined below.
@@
We perform several additional steps to ensure high-quality mistake detection task instances. First, we remove clips where the precondition and postcondition frames are overly similar (i.e., at least 0.95 cosine similarity). We remove clips that are too dark (i.e., where the mean of all normalized RGB values is less than 0.2). When sampling frames from source clips, we sample several candidates within a small range around the precondition or postcondition timestamp, then select the least blurry candidate by the variance of the images’ Laplacian. Some videos in Ego4D show the same action being performed over and over (e.g., “Roll a ball of dough”), which can make it difficult to determine whether the state of the environment shown in a clip is the result of the current procedure or a prior one (given only a single frame). While future work applying video-optimized VLMs for explainable PMD in long-horizon tasks will need to address this challenge, this adds an unnecessary complexity to an already challenging task for current VLMs. As such, we remove any clips such that the same procedure in the clip has already been performed previously in the video. Lastly, we remove a handful of videos in Ego4D that we notice to be corrupted or significantly distorted.
Various statistics of the resulting full Ego4D-PMD dataset (and sub-samples used in the paper) are presented in Tables 3 and 4. It is worth noting that this collected dataset does not include ground truth explanations for mistake detection labels (rather we proposed automated reference-free metrics) for two important reasons. First, there may be multiple valid ways to detect a mistake through asking and answering visual questions, each of which could consist of different questions and/or different numbers of questions. [^17] For example, in trying to determine the success of the procedure “In a bowl, add the cut cherry tomatoes” from Peddi et al. (2024), we could reasonably ask one question “Are all the cherry tomatoes in the bowl?” or two questions “Are there cherry tomatoes in the bowl?” and “Are there any cherry tomatoes outside of the bowl?” In our opinion, existing metrics for text generation are not as well suited for this extremely challenging evaluation as fine-tuned NLI models, which are optimized to judge the logical consistency between texts. Second, in a real-world setting like PMD for task guidance, we believe that automated metrics are better suited for understanding and improving a deployed system than an offline benchmark of ground truth explanations.
@@ Appendix C Coherence Metrics Human Study Details
In this appendix, we collect human judgments for the relevance and informativeness metrics defined in Section 3. To achieve this, we randomly sampled the outputs for 50 iterations of LLaVA’s self-dialog from two combinations of evaluated approaches:
@@ C.1 Annotation Instructions
We recruit 5 annotators, each of which annotates 10 samples. One sample is listed below:
@@
We recruit 5 annotators, each of which annotates 10 samples. One sample is listed below:
@@ C.2 Human Study Results
For each metric, we recruited 5 annotators (all English speakers with with conferred or in-progress undergraduate degrees) to rate the relevance and informativeness of machine-generated questions and answers.
@@
We presented each annotator with 10 randomly selected VLM-generated questions $Q^{\prime}$, along with previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$. Annotators were instructed to rate the relevance (i.e., given the previous questions and answers, how helpful could an answer to this question be in determining whether the task was successfully completed) on a scale from 1-5 (least to most relevant). Between the resulting 50 annotations and corresponding automated metrics, we found a moderate Spearman correlation Spearman (1904) of $\rho=0.55$ ($p=0.000029$). This suggests that this automated measure of relevance is indeed positively correlated with human judgments of relevance.
Comparison of automated and human informativeness judgments.
We presented each annotator with 10 randomly selected VLM-generated questions $Q^{\prime}$ and answers $A^{\prime}$, along with previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$. Annotators were instructed to rate the relevance (i.e., based on all the information we have, how sure is the annotator about whether the procedure was successfully completed) on a scale from 1-5 (least to most informative). Between the resulting 50 annotations and corresponding automated metrics, we found a weaker Spearman correlation Spearman (1904) of $\rho=0.32$ ($p=0.024$). Interestingly, if we multiply the automated informativeness metrics by the relevance for $Q^{\prime}$, $\mathcal{Q}$, and $\mathcal{A}$, we find a stronger Spearman correlation of $\rho=0.50$ ($p=0.00022$). This suggests that while informativeness does have a relationship with human judgments, when multiplying it by relevance this relationship is stronger and more significant. This might be because the concepts of relevance and informativeness are themselves related. Intuitively, in most cases, a relevant question should be informative, and an irrelevant question should be uninformative. Meanwhile, fine-tuned NLI models could theoretically score an answer to an irrelevant question as informative, and vice versa. When proposing a coherence-based question re-ranking strategy in Section 4.2.1, we incorporated the inductive bias reflected in these human judgments by multiplying informativeness by relevance. Later, in Section 4.4, we visualized the full distribution of relevance and informativeness on our evaluation data to better understand this issue.
@@ Appendix D VLM Self-Dialog Details
In this appendix, we include prompt templates and other supplementary details for the self-dialog we conditioned VLMs to generate, including visual question generation, visual question answering, and stopping criteria. To conserve GPU memory, VLM weights are 4-bit quantized at inference time.
@@ D.1 Visual Question Generation
Question generation is not conditioned on video frames, as we found significant performance degradation when VLMs were conditioned on the video frame while generating questions, often leading to completely nonsensical questions, e.g., “Is is is is is is?” To ensure the VLM generates yes-no questions, we constrain generation during decoding to enforce that each generated text begins with a word that can signal a yes-no question, [^18] Specifically, questions must begin with is, does, or has, along with all plural and past tense forms of these verbs. does not include the word or, and ends with a question mark. To avoid vague and high-level questions about the status of the procedure rather than low-level physical states, we prevent VLMs from repeating some words present in their input prompts: “successful,” “successfully,” “executed,” and “procedure.” To encourage logical questions while ensuring variety, we apply greedy beam search decoding with $k=8$ beams, returning the top 4 candidate questions to be ranked through LM likelihood or coherence metrics. [^19] Due to generation constraints, it is often the case that the VLM does not successfully generate all 8 candidates. In rare cases, these constraints prevent VLMs from generating any candidate questions; when this happens, we repeat the generation without any constraints (even though this may result in outputs that are not yes-no questions). Out of the candidates, we remove any that are exactly the same as previously generated questions, then select the most likely candidate based on the model’s log-likelihoods.
@@ D.4 Stopping Criteria Details
These hyperparameters $n$, $\delta$, and $\epsilon$ for the self-dialog stopping crtieria are selected from a grid search on the validation data over combinations of $n\in\{2,4,6,8,10\}$, $\delta\in\{0.05,0.1,0.2,0.4\}$, and $\epsilon\in\{0.025,0.05,0.1,0.2\}$. We maximize overall accuracy, relevance of questions, and potential informativeness of their answers. Specifically, we use a summary metric which is the product of example-level informativeness (reference-adjusted) and relevance if and only if the VLM makes an accurate mistake detection prediction and this product is positive, else it is zero. Accuracy depends on the mistake confidence threshold $\tau$, which is selected as described above. Selected stopping criteria hyperparameter values for the results in Table 2 are listed later in Appendix E.2, while values for the results in Table 2 are listed in Appendix E.5.
Appendix E Experiment Details and Supplementary Results
In this appendix, we provide assorted details (e.g., hyperparameters) and supplementary results that were omitted from the main body of the paper.
E.1 Example Questions for In-Context Learning in Question Generation
As discussed in Section 4.2.2, we condition VLMs with sets of human-written questions for 20 procedures from the Ego4D for Procedural Mistake Detection (Ego4D-PMD) dataset. The annotated procedures (underlined) and questions (italicized) are listed below:
@@ E.1 Example Questions for In-Context Learning in Question Generation
E.2 Question Selection Inference Hyperparameters and Validation Results
@@
E.3 Analysis of Question Sources in In-Context Learning
To shed more light on where selected candidate questions come from in each approach, we visualize the distribution of question sources on the validation data in Figure 6. As expected, candidates generated with in-context learning are only rarely selected in the likelihood-based ranking, amounting to about 26.3% of VQG iterations for InstructBLIP, 9.0% of VQG iterations for LLaVA, and 4.9% of VQG iterations for Llama 3. On the other hand, they are selected more frequently in the coherence-based ranking, amounting to about 35.3% of VQG iterations for InstructBLIP, 38.0% of VQG iterations for LLaVA, and 36.5% of VQG iterations for Llama 3. Interestingly, in-context learning candidates are more dominant in earlier iterations, while candidates generated based on the dialog context are relatively more common in later iterations. This may suggest that after selecting a few questions from in-context learning in earlier iterations, the VLM is able to utilize them to generate better questions from the dialog context in later iterations. Alternatively, this could suggest that candidates from in-context learning have limited variety, and thus are less likely to be selected in later turns to avoid redundant questions or information.

### v2

@@ Acknowledgments
This work was supported in part by the DARPA PTG program HR00112220003. We would like to thank Megan Su, Ruixuan Deng, Fengyuan Hu, Andy Chung, Lu Wang, Richard L. Lewis, and the entire MSRP team for their helpful discussions, feedback, and insights.
@@ Appendix A Ego4D-PMD Data Curation Details
To alleviate these challenges and focus our inquiry, we follow Du et al. (2023) in recasting Ego4D, a procedural video dataset with a breadth of annotated and narrated everyday actions, into a single frame mistake detection format. The collection of the Ego4D Grauman et al. (2022) for Procedural Mistake Detection (Ego4D-PMD) data consisted of several preprocessing steps, outlined below.
@@
We perform several additional steps to ensure high-quality mistake detection task instances. First, we remove clips where the precondition and postcondition frames are overly similar (i.e., at least 0.95 cosine similarity). We remove clips that are too dark (i.e., where the mean of all normalized RGB values is less than 0.2). When sampling frames from source clips, we sample several candidates within a small range around the precondition or postcondition timestamp, then select the least blurry candidate by the variance of the images’ Laplacian. Some videos in Ego4D show the same action being performed over and over (e.g., “Roll a ball of dough”), which can make it difficult to determine whether the state of the environment shown in a clip is the result of the current procedure or a prior one (given only a single frame). While future work applying video-optimized VLMs for coherent PMD in long-horizon tasks will need to address this challenge, this adds an unnecessary complexity to an already challenging task for current VLMs. As such, we remove any clips such that the same procedure in the clip has already been performed previously in the video. Lastly, we remove a handful of videos in Ego4D that we notice to be corrupted or significantly distorted.
Various statistics of the resulting full Ego4D-PMD dataset (and sub-samples used in the paper) are presented in Tables 3 and 4.
@@ Appendix C Coherence Metrics Human Study Details
In this appendix, we collect human judgments for the relevance and informativeness metrics defined in Section 3. To achieve this, we randomly sampled the outputs for 50 iterations of LLaVA’s self-dialog from two combinations of approaches:
@@
These outputs were from intermediate experiments, and thus the prompt used to initiate the self-dialog is slightly different than the one used for the experiments in the paper. However, as this is just a source of questions to compare human and machine judgments, this discrepancy does not impact the conclusions in the paper.
@@ C.1 Annotation Instructions
One sample is listed below:
@@
One sample is listed below:
@@ C.2 Human Study Results
For each metric, we recruited 5 annotators (all English speakers with conferred or in-progress undergraduate degrees) to rate the relevance and informativeness of 10 machine-generated questions and answers. As this was an initial pilot study that was not scaled up further, annotators were authors of the work (not the first author) and/or peers of the authors, and not compensated. When comparing with our automated relevance and informativeness metrics, we use LLaVA for rephrasing.
Comparison of automated and human relevance judgments.
We presented each annotator with 10 randomly selected VLM-generated questions $Q^{\prime}$, along with previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$. Annotators were instructed to rate the relevance (i.e., given the previous questions and answers, how helpful could an answer to this question be in determining whether the task was successfully completed) on a scale from 1-5 (least to most relevant). Between the resulting 50 annotations and corresponding automated metrics, we found a moderate Spearman correlation Spearman (1904) of $\rho=0.55$ ($p=0.000037$). This suggests that this automated measure of relevance is indeed positively correlated with human judgments of relevance.
@@
We presented each annotator with 10 randomly selected VLM-generated questions $Q^{\prime}$ and answers $A^{\prime}$, along with previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$. Annotators were instructed to rate the relevance (i.e., based on all the information we have, how sure is the annotator about whether the procedure was successfully completed) on a scale from 1-5 (least to most informative). Between the resulting 50 annotations and corresponding automated metrics, we found a weaker Spearman correlation Spearman (1904) of $\rho=0.33$ ($p=0.020$). Interestingly, if we multiply the automated informativeness metrics by the relevance for $Q^{\prime}$, $\mathcal{Q}$, and $\mathcal{A}$, we find a stronger Spearman correlation of $\rho=0.50$ ($p=0.00022$). This suggests that while informativeness does have a relationship with human judgments, when multiplying it by relevance this relationship is stronger and more significant. This might be because the concepts of relevance and informativeness are themselves related. Intuitively, in most cases, a relevant question should be informative, and an irrelevant question should be uninformative. Meanwhile, fine-tuned NLI models could theoretically score an answer to an irrelevant question as informative, and vice versa. When proposing a coherence-based question re-ranking strategy in Section 4.1.1, we incorporated the inductive bias reflected in these human judgments by multiplying informativeness by relevance. Later, in Section 5.3, we visualized the full distribution of relevance and informativeness on our evaluation data to better understand this issue.
@@ Appendix D VLM Self-Dialog Details
In this appendix, we include prompt templates and other supplementary details for the self-dialog we conditioned VLMs to generate, including visual question generation, visual question answering, and stopping criteria. To conserve GPU memory, VLM weights are 4-bit quantized at inference time, and are obtained from https://huggingface.co/Salesforce/instructblip-vicuna-7b, https://huggingface.co/liuhaotian/llava-v1.5-7b, and https://huggingface.co/meta-llama/Llama-3.2-11B-Vision.
@@ D.1 Visual Question Generation
Question generation is not conditioned on video frames, as we found significant performance degradation when VLMs were conditioned on the video frame while generating questions, often leading to completely nonsensical questions, e.g., “Is is is is is is?” To ensure the VLM generates yes-no questions, we constrain generation during decoding to enforce that each generated text begins with a word that can signal a yes-no question, [^21] Specifically, questions must begin with is, does, or has, along with all plural and past tense forms of these verbs. does not include the word or, and ends with a question mark. To avoid vague and high-level questions about the status of the procedure rather than low-level physical states, we prevent VLMs from repeating some words present in their input prompts: “successful,” “successfully,” “completed,” and “procedure.” To encourage logical questions while ensuring variety, we apply greedy beam search decoding with $k=8$ beams, returning the top 4 candidate questions to be ranked through LM likelihood or coherence metrics. [^22] Due to generation constraints, it is often the case that the VLM does not successfully generate all 8 candidates. In rare cases, these constraints prevent VLMs from generating any candidate questions; when this happens, we repeat the generation without any constraints (even though this may result in outputs that are not yes-no questions). Out of the candidates, we remove any that are exactly the same as previously generated questions, then select the most likely candidate based on the model’s log-likelihoods.
@@ D.4 Stopping Criteria Details
The stopping criteria hyperparameter $n$ is fixed at a value of 10. We chose to fix $n$ because tuning $n$ prevents objective comparison of the number of iterations taken by each approach, and $\delta$ and $\epsilon$ can still control the number of iterations of VLMs’ self-dialog. The hyperparameters $\delta$ and $\epsilon$ are selected from a grid search on the validation data over combinations of $\delta\in\{0.05,0.1,0.2,0.4\}$ and $\epsilon\in\{0.025,0.05,0.1,0.2\}$. We maximize overall accuracy, relevance of questions, and potential informativeness of their answers. Specifically, we use a cascading summary metric which is the product of example-level informativeness (reference-adjusted) and example-level relevance if the VLM makes a correct mistake detection classification and this product is positive, else it is zero. Accuracy depends on the mistake confidence threshold $\tau$, which is selected as described in the previous subsection. Selected stopping criteria hyperparameter values for the results in Table 2 are listed later in Appendix E.6, while values for the results in Table 2 are listed in Appendix E.3.
Appendix E Supplementary Experimental Details and Results
In this appendix, we provide assorted details (e.g., prompt templates and hyperparameters) and supplementary results that were omitted from the main body of the paper.
E.1 Example Questions for In-Context Learning
As discussed in Section 4.1.2, we condition VLMs with sets of human-written questions for 20 procedures from the Ego4D for Procedural Mistake Detection (Ego4D-PMD) dataset. These human-written questions individually achieve 53.2% relevance and 82.1% maximum informativeness (i.e., for a yes or no answer) on average, with rephrasing done by LLaVA. The annotated procedures (underlined) and questions (italicized) are listed below:
@@ E.1 Example Questions for In-Context Learning
E.2 DPO Training Data Composition Ablation
In fine-tuning VLMs to generate questions, we included questions generated through the in-context learning approach introduced in Section 4.1.2 to build upon our previous best result (i.e., augmenting VLMs with coherence-based ranking and in-context learning). In Table 5, we include an additional result for DPO where generated training data does not include candidate questions generated with in-context learning. As shown, compared to the results in Table 2, the best values of each metric are largely comparable with or without in-context learning, suggesting that in-context learning did not contribute significantly to the performance.
LLaVA + DPO (without ICL)
L 61.0 84.3 .312 2.16 .746
L 64.8 64.9 .326 2.76 .622
C 61.2 94.6 .316 1.78 .772
C 61.8 93.0 .342 2.37 .790
E.3 Question Generation Fine-Tuning and Inference Details
[…]
LLaVA + DPO
$n$ $\delta$ $\epsilon$
10 0.05 0.05
LLaVA + DPO (without ICL)
$n$ $\delta$ $\epsilon$
10 0.05 0.05
LLaVA + DPO (with length penalty)
$n$ $\delta$ $\epsilon$
10 0.05 0.05
Table 6: Selected training data generation hyperparameters for the results presented in Tables 2, 5, and 17.
LLaVA + DPO
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
L 1e-5 0.5 10 0.2 0.05 0.78
L 10 0.2 0.05 0.75
C 10 0.2 0.05 0.58
C 10 0.1 0.1 0.76
LLaVA + DPO (without ICL)
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
L 7.5e-6 0.5 10 0.1 0.05 0.59
L 10 0.1 0.05 0.85
C 10 0.4 0.05 0.66
C 10 0.05 0.05 0.64
LLaVA + DPO (with length penalty)
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
L 7.5e-6 0.1 10 0.4 0.05 0.35
L 10 0.05 0.05 0.63
C 10 0.4 0.05 0.86
C 10 0.4 0.05 0.84
Table 7: Selected training and inference hyperparameters for the results presented in Tables 2, 5, and 17.
LLaVA + DPO
L 63.0 76.7 .328 2.27 .617
L 64.0 65.5 .347 2.48 .579
C 63.0 91.2 .342 2.11 .703
C 65.2 94.8 .318 1.85 .727
LLaVA + DPO (without ICL)
L 62.0 85.0 .293 2.04 .757
L 62.6 71.9 .291 2.40 .689
C 62.6 93.6 .338 1.85 .756
C 63.0 92.0 .357 2.54 .783
LLaVA + DPO (with length penalty)
L 62.0 93.1 .293 1.97 .739
L 65.4 66.1 .412 3.61 .648
C 61.6 97.1 .281 1.86 .756
C 61.8 97.2 .290 1.85 .758
Table 8: Ego4D-PMD validation set results for DPO-trained VLMs both with and without applying in-context learning in generating training data, as well as with a length penalty applied in generating training data. Inference applies likelihood (L) or coherence (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.4 GPT-4o Evaluation
To better contextualize our results with state-of-the-art proprietary large LMs, we additionally evaluated GPT-4o OpenAI et al. (2024b) [^23] Specifically, we use the August 6, 2024 version of GPT-4o available in Azure AI Foundry. on our Ego4D-PMD dataset for coherent PMD. While proprietary models like GPT-4o offer limited customization, making many of the experiments we presented for open-source VLMs, evaluating vanilla GPT-4o serves as a reference point. It is worth noting that even accessed through APIs, GPT-4o returns responses too slowly to be viable for online, frame-by-frame use.
This is…from the photo. Generate an appropriate yes/no question.
Q:
We also ran into a couple issues with the Azure OpenAI API. Some specific examples in our dataset triggered Azure’s content filter, even when setting the filter’s threshold to high. In the cases where the issue persisted, we were forced to skip over the example during evaluation. Also, a small portion of API responses returned no completion content during VQG or question rephrasing for NLI (i.e. the content attribute of the response had a value of None). To deal with this, we implemented a retry procedure, where if no content is given by GPT, we prompt it again with the same prompt. If this second chance also provides no content, we either ignore the example during evaluation in the case of VQG, or simply concatenate the question and answer for question rephrasing. On the validation data, 28 out of 500 examples were omitted, and 103 question candidates could not be rephrased by GPT-4o.
The inference hyperparameters $n$, $\delta$, $\epsilon$, and $\tau$ are selected as in our open-source model results, and listed in Table 9. The results on the validation data are listed in Table 10, while the results on the testing data are listed in the main body of the paper in Table 2. Comparing GPT-4o to the base VLMs we evaluated, it is generally inferior in PMD accuracy and informativeness, but asks more relevant questions, runs for fewer iterations, and has higher information gain. Our best model configurations, though, outperform GPT-4o under all evaluation metrics except information gain. [^24] LLaVA with coherence-based fine-tuning, coherence-based re-ranking, and in-context learning achieves .742 bits of information gain, while GPT-4o achieves up to .793 bits. This shows that GPT-4o generally exhibits higher confidence despite having much lower PMD accuracy than our approaches, which is not necessarily an advantage. This suggests that GPT-4o is a reasonable starting point for coherent PMD, but like other off-the-shelf VLMs we evaluated, it may require additional interventions (e.g., coherence-based ranking) to be viable for this task.
$n$ $\delta$ $\epsilon$ $\tau$
10 0.4 0.1 0.19
Table 9: Selected inference hyperparameters for GPT-4o.
Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
58.7 54.3 .175 1.83 .730
Table 10: Ego4D-PMD validation set results for GPT-4o.
E.5 Rationale-Free Evaluation
For a reference point to incomparable prior works that have applied VLMs to PMD with a focus on classification accuracy Du et al. (2023); Bao et al. (2023), we additionally evaluate all studied VLMs from this work on a non-coherent PMD task. Specifically, we prompt each VLM with the following text:
[…]
This prompt is as comparable as possible to the one used for coherent PMD, but does not elicit a series of questions and answers from the VLM. We perform this evaluation on InstructBLIP, LLaVA, Llama 3, and GPT-4o. Table 11 lists the inference hyperparameters for this approach, while Table 12 lists the results for the validation and testing data. It is crucial to note that these results are not directly comparable to the coherent PMD results presented in the main body of the paper, as VLMs are not required to justify their decisions, removing the explainability enabled by coherent PMD (crucial for end users to interpret often incorrect VLM decisions and act on them accordingly). As prior work has already explored this setting more extensively, we do not intend to provide a rigorous study here, rather a reference point to compare how requiring the generation of rationales impacts PMD accuracy.
Interestingly, however, the rationale-free approach achieves generally better accuracy than the coherent PMD results in Tables 2 and 14 with likelihood-based ranking and no in-context learning. However, upon introducing coherence-based ranking and in-context learning, InstructBLIP and LLaVA achieve better accuracy than in the rationale-free approach. Further, the information gain in the rationale-free setting is consistently lower than those achieved in coherent PMD. This demonstrates that while the added transparency of incorporating rationales into PMD does cost the VLM some accuracy, improving the coherence of these rationales (e.g., through the approaches presented in this work) can recover this accuracy and more while enabling more confident decisions from VLMs. Rationale-free GPT-4o achieves the highest observed accuracy of 69.2%. Nonetheless, this accuracy is still low enough for errors to be common, thus necessitating the generation of a rationale for the user.
InstructBLIP LLaVA Llama 3 GPT-4o
0.34 0.77 0.54 0.98
Table 11: Selected values of mistake confidence threshold $\tau$ for rationale-free PMD with various VLMs. $n$, $\delta$, and $\epsilon$ are no longer used in rationale-free PMD, but $\tau$ is still tuned as in previous experiments.
InstructBLIP
Validation 62.6 0.00 .113
Testing 62.2 0.00 .117
LLaVA
Validation 64.4 0.00 .236
Testing 66.1 0.00 .233
Llama 3
Validation 65.8 0.00 .187
Testing 64.6 0.00 .176
GPT-4o
Validation 66.8 0.00 .718
Testing 69.2 0.00 .736
Table 12: Ego4D-PMD validation and testing set results for rationale-free PMD with various VLMs. As self-dialog rationales are no longer generated, relevance and informativeness cannot be calculated. Further, zero iterations are performed, and information gain is calculated using the average entropy of the success probability without any rationale.
E.6 Question Selection Inference Hyperparameters and Validation Results
@@
E.7 Analysis of Question Sources in In-Context Learning
To shed more light on where selected candidate questions come from in each approach, we visualize the distribution of question sources on the validation data in Figure 6. As expected, candidates generated with in-context learning are relatively rarely selected under likelihood-based ranking, amounting to about 25.7% of VQG iterations for InstructBLIP, 9.0% of VQG iterations for LLaVA, and 5.2% of VQG iterations for Llama 3. On the other hand, they are selected more frequently in the coherence-based ranking, amounting to about 35.6% of VQG iterations for InstructBLIP, 30.6% of VQG iterations for LLaVA, and 36.5% of VQG iterations for Llama 3. Interestingly, in-context learning candidates are more dominant in earlier iterations, while candidates generated based on the dialog context are relatively more common in later iterations. This may suggest that after selecting a few questions from in-context learning in earlier iterations, the VLM is able to utilize them to generate better questions from the dialog context in later iterations. Alternatively, this could suggest that candidates from in-context learning have limited variety, and thus are less likely to be selected in later turns to avoid redundant questions or information.